# MEDebiaser: A Human-AI Feedback System for Mitigating Bias in Multi-label Medical Image Classification

### v2

@@ 5.2.4. Result Processing and Measurement
For the post-task questionnaire, a 7-point Likert scale was employed ($1$: Not at all/Strongly disagree, $7$: Very much/Strongly agree). Further details are available in Table 7. The questions were categorized into four aspects: bias detection, system usability scale (Brooke, 2013), feedback & communication, and workload (Hart and Staveland, 1988). Among these, the system usability scale and bias detection primarily addressed RQ4, while feedback & communication and workload primarily addressed RQ5. Even though engineers do not directly interact with the system, their feedback on bias detection, usability, and workload can provide valuable insights into the system’s overall performance in a collaborative setting. Thus, as observers, they also filled out the same questionnaire.
@@ 5.3. User Study Results
We compiled information on users’ interactions with MEDebiaser through the study. Initially, all physicians successfully uploaded their datasets and selected a model in the Panel View. They then proceeded to the Label View, where five physicians sorted the table in ascending order by distribution percentage. These physicians selected categories with lower distribution, such as EOM, OME, and EOF, and closely examined the label co-occurrence in the chord diagram. In the Attention View, they reviewed $38$ images with a total of $49$ labels under EOF. Among these images, $18$ were correctly predicted, covering $25$ labels, while $20$ images were incorrectly predicted. Four physicians zoomed in on each image to correct biases in unreasonable predictions, with UD3 and UD4 opting to directly select correctly predicted attention and adjust only for incorrect predictions. In the Modification View, all six physicians adeptly used polygons to annotate the images, and UD1 and UD5 also utilized the dropdown menu to annotate the remaining labels. For recommendation, UD1, UD2, and UD6 chose the dependency mode for sorting, UD3 and UD5 selected concentration, and UD4 opted for accuracy. In the Performance View, the physicians reviewed EOF, carefully comparing each image with the results from the previous round. During this stage, the physicians discussed the model’s metrics and collaborated with engineers to get further assurance.
@@
Regarding imbalance mitigation (physicians: mean = $5.67$; engineers: mean = $4.83$) and co-occurrence distinction (physicians: mean = $5.17$; engineers: mean = $4.5$), it is clear from Figure 5 that physicians rated these aspects higher than engineers. Several physicians observed that, in the Performance View, after multiple rounds of fine-tuning, the attention seems to align more accurately with the correct regions. However, two engineers mentioned that the relatively small amount of annotated data in the user study has led to only minor changes in metrics. UP1 commented, “Right now, with the limited number of annotations, the metrics only show small improvements. But I think as we add more [annotations], we’ll start to see a bigger reduction [in bias].
@@
MEDebiaser is easy to use, and its interactive human-AI feedback mechanism effectively aids users in correcting model biases with ease. Both physicians and engineers acknowledged the ease of use of MEDebiaser (physicians: mean = $6$; engineers: mean = $6.33$) and its low learning curve (physicians: mean = $6.33$; engineers: mean = $5.83$), as shown in Figure 5. Notably, UD2, UD5, and UD6, despite having no prior ML experience, also found MEDebiaser accessible and agreed that it requires minimal ML knowledge to operate. UD3 mentioned, “MEDebiaser is quite user-friendly and easy to use. It’s similar to Labelme (Russell et al., 2008), which I have used before for image annotation.” As shown in Table 8, all six teams completed the modification work for EOF within the designated time (T3). Among them, UD4 completed the task with only three rounds of fine-tuning in the shortest time, while UD5 took the longest, finishing in five rounds. This difference is attributed to UD4’s preference for creating rough outlines with fewer points, compared to UD5’s more detailed approach with precise points, as shown in Figure 6. This iterative feedback mechanism allowed physicians to observe changes and trends in the model, facilitating dynamic adjustments and enhancing efficiency. As UD1 noted, “After two rounds of adjustments, I could see the results getting better (Figure 7). I was thinking of moving on to other [labels], but I ended up sticking with the remaining [EOF] images as needed.” Overall, both physicians and engineers expressed a willingness to use MEDebiaser in the future (physicians: mean = $5$; engineers: mean = $5.33$).
Figure 6. 1 Original Image. 2 Polygon annotated by UD5. 3 Polygon annotated by UD4.
@@
Figure 7. The results of two rounds of fine-tuning on EOF by UD1. In 1, the red box contains a black foreign object. After each fine-tuning, the attention gets closer to the foreign object. After two rounds of fine-tuning, the attention covers the foreign object. In 2, the red box contains a strand of hair. After each fine-tuning, the attention gets closer to the hair. After two rounds of fine-tuning, the attention covers the entire hair.
@@
MEDebiaser’s interface effectively facilitates iterative feedback between physicians and AI models, promoting the integration of expertise into the training process. MEDebiaser employs Grad-CAM to provide local explanations. Despite being a conventional method, both physicians and engineers found this visualization straightforward and easy to interpret (physicians: mean = $5.67$; engineers: mean = $5.83$), as shown in Figure 8. UD6 pointed out, “We frequently request visualizations, but sometimes they end up being overly complicated and hard to use. Most of the time, simple heatmaps do the job perfectly well.” Regarding the integration of medical expertise, physicians expressed high satisfaction (mean = $6.33$). UD1 mentioned, “Previously, it was really challenging for me to communicate the important [medical] features I had in mind to engineers and have them implemented. However, with this new approach, my ideas are now directly and accurately represented in the model, and the heatmaps and line charts it produces are clear and easy to understand.” (T2, T4). Engineers found this method of knowledge integration more accessible to physicians and less complex compared to existing approaches (mean = $5.67$). UP3 commented, “MEDebiaser’s interface is designed with clarity, making it easy to interpret the model and guide users effectively. We often use iterative training methods in our work, but usually lack a user interface to facilitate these [processes].
@@
MEDebiaser has effectively reduced both the time and physical workload for physicians and engineers. Most physicians found its interactive mechanism significantly decreased the number of required annotations compared to annotating the entire dataset before training, reducing time (mean = $5.67$) and physical workload (mean = $5.33$). However, UD2 commented, “Even though it cuts down on the number of annotations compared to labeling the entire dataset, I still don’t want to spend too much [time] on this annotation [work].” Regarding the physical workload, physicians had varying opinions. UD4, who used fewer points for each annotation, felt the physical workload was manageable, while UD5, who meticulously annotated each image, expressed concern: “When there’s a lot of data, this work can get exhausting, both physically and mentally.” engineers also reported reduced time workload (mean = $4.83$) and physical workload (mean = $4.67$). UP1 noted, “Our main role is to guide the training settings and confirm the results, which is a lot easier than having to fix the model whenever something goes wrong, especially when we don’t fully understand the data.
@@ Appendix B Loading Dataset and Model Supplementary
  • Atl = Atelectasis, Car = Cardiomegaly, Con = Consolidation, Ede = Edema, Eff = Effusion, Emp = Emphysema, Fib = Fibrosis, Her = Hernia, Inf = Infiltration, Mas = Mass, Nod = Nodule, Ple = Pleural_Thickening, Pne = Pneumonia, Ptx = Pneumothorax
@@ Appendix C Observing and Modifying Attention Supplementary
Figure 10. The comparison of different visualization and interpretability methods is as follows: 1 Original Chest X-ray: Provide the baseline image for analysis. 2 Grad-CAM: Highlights the most influential regions in a coherent, high-level manner. 3 LIME (Top 10 Regions): Delivers fragmented, isolated areas that can be difficult to interpret within a broader diagnostic context. 4 Guided Backpropagation: May introduce noise, potentially obscuring clinically relevant features, particularly in multi-label settings. 5 SHAP: Often results in dispersed or sparse patterns, making interpretation more challenging.
@@ Appendix D Evaluation Metrics Supplementary
where $R_{i}$ is the set of retrieved results for the $i$-th class, ordered by their confidence scores; $P(r)$ is the precision at the $r$-th rank; $\Delta r$ is the change in recall from the $(r-1)$-th to the $r$-th retrieved result.
@@ Appendix E User Study Supplementary
  • OME = Otitis Media with Effusion, EOF = External Auditory Canal Foreign Body, OE = Otitis Externa, EOM = External Auditory Canal Tumor, FOE = Fungal Otitis Externa, CE = Cerumen Impaction, TMP = Tympanic Membrane Perforation

### v3

@@ 5.2.4. Result Processing and Measurement
For the post-task questionnaire, a 7-point Likert scale was employed (111: Not at all/Strongly disagree, 777: Very much/Strongly agree). Further details are available in Table 7. The questions were categorized into four aspects: bias detection, system usability scale (Brooke, 2013), feedback & communication, and workload (Hart and Staveland, 1988). Among these, the system usability scale and bias detection primarily addressed RQ4, while feedback & communication and workload primarily addressed RQ5. Even though engineers do not directly interact with the system, their feedback on bias detection, usability, and workload can provide valuable insights into the system’s overall performance in a collaborative setting. Thus, as observers, they also filled out the same questionnaire.
@@ 5.3. User Study Results
We compiled information on users’ interactions with MEDebiaser through the study. Initially, all physicians successfully uploaded their datasets and selected a model in the Panel View. They then proceeded to the Label View, where five physicians sorted the table in ascending order by distribution percentage. These physicians selected categories with lower distribution, such as EOM, OME, and EOF, and closely examined the label co-occurrence in the chord diagram. In the Attention View, they reviewed 383838 images with a total of 494949 labels under EOF. Among these images, 181818 were correctly predicted, covering 252525 labels, while 202020 images were incorrectly predicted. Four physicians zoomed in on each image to correct biases in unreasonable predictions, with UD3 and UD4 opting to directly select correctly predicted attention and adjust only for incorrect predictions. In the Modification View, all six physicians adeptly used polygons to annotate the images, and UD1 and UD5 also utilized the dropdown menu to annotate the remaining labels. For recommendation, UD1, UD2, and UD6 chose the dependency mode for sorting, UD3 and UD5 selected concentration, and UD4 opted for accuracy. In the Performance View, the physicians reviewed EOF, carefully comparing each image with the results from the previous round. During this stage, the physicians discussed the model’s metrics and collaborated with engineers to get further assurance.
@@
Regarding imbalance mitigation (physicians: mean = 5.675.675.67; engineers: mean = 4.834.834.83) and co-occurrence distinction (physicians: mean = 5.175.175.17; engineers: mean = 4.54.54.5), it is clear from Figure 5 that physicians rated these aspects higher than engineers. Several physicians observed that, in the Performance View, after multiple rounds of fine-tuning, the attention seems to align more accurately with the correct regions. However, two engineers mentioned that the relatively small amount of annotated data in the user study has led to only minor changes in metrics. UP1 commented, “Right now, with the limited number of annotations, the metrics only show small improvements. But I think as we add more [annotations], we’ll start to see a bigger reduction [in bias].
@@
MEDebiaser is easy to use, and its interactive human-AI feedback mechanism effectively aids users in correcting model biases with ease. Both physicians and engineers acknowledged the ease of use of MEDebiaser (physicians: mean = 666; engineers: mean = 6.336.336.33) and its low learning curve (physicians: mean = 6.336.336.33; engineers: mean = 5.835.835.83), as shown in Figure 5. Notably, UD2, UD5, and UD6, despite having no prior ML experience, also found MEDebiaser accessible and agreed that it requires minimal ML knowledge to operate. UD3 mentioned, “MEDebiaser is quite user-friendly and easy to use. It’s similar to Labelme (Russell et al., 2008), which I have used before for image annotation.” As shown in Table 8, all six teams completed the modification work for EOF within the designated time (T3). Among them, UD4 completed the task with only three rounds of fine-tuning in the shortest time, while UD5 took the longest, finishing in five rounds. This difference is attributed to UD4’s preference for creating rough outlines with fewer points, compared to UD5’s more detailed approach with precise points, as shown in Figure 6. This iterative feedback mechanism allowed physicians to observe changes and trends in the model, facilitating dynamic adjustments and enhancing efficiency. As UD1 noted, “After two rounds of adjustments, I could see the results getting better (Figure 7). I was thinking of moving on to other [labels], but I ended up sticking with the remaining [EOF] images as needed.” Overall, both physicians and engineers expressed a willingness to use MEDebiaser in the future (physicians: mean = 555; engineers: mean = 5.335.335.33).
Figure 6. Refer to caption Original Image. Refer to caption Polygon annotated by UD5. Refer to caption Polygon annotated by UD4.
@@
Figure 7. The results of two rounds of fine-tuning on EOF by UD1. In Refer to caption, the red box contains a black foreign object. After each fine-tuning, the attention gets closer to the foreign object. After two rounds of fine-tuning, the attention covers the foreign object. In Refer to caption, the red box contains a strand of hair. After each fine-tuning, the attention gets closer to the hair. After two rounds of fine-tuning, the attention covers the entire hair.
@@
MEDebiaser’s interface effectively facilitates iterative feedback between physicians and AI models, promoting the integration of expertise into the training process. MEDebiaser employs Grad-CAM to provide local explanations. Despite being a conventional method, both physicians and engineers found this visualization straightforward and easy to interpret (physicians: mean = 5.675.675.67; engineers: mean = 5.835.835.83), as shown in Figure 8. UD6 pointed out, “We frequently request visualizations, but sometimes they end up being overly complicated and hard to use. Most of the time, simple heatmaps do the job perfectly well.” Regarding the integration of medical expertise, physicians expressed high satisfaction (mean = 6.336.336.33). UD1 mentioned, “Previously, it was really challenging for me to communicate the important [medical] features I had in mind to engineers and have them implemented. However, with this new approach, my ideas are now directly and accurately represented in the model, and the heatmaps and line charts it produces are clear and easy to understand.” (T2, T4). Engineers found this method of knowledge integration more accessible to physicians and less complex compared to existing approaches (mean = 5.675.675.67). UP3 commented, “MEDebiaser’s interface is designed with clarity, making it easy to interpret the model and guide users effectively. We often use iterative training methods in our work, but usually lack a user interface to facilitate these [processes].
@@
MEDebiaser has effectively reduced both the time and physical workload for physicians and engineers. Most physicians found its interactive mechanism significantly decreased the number of required annotations compared to annotating the entire dataset before training, reducing time (mean = 5.675.675.67) and physical workload (mean = 5.335.335.33). However, UD2 commented, “Even though it cuts down on the number of annotations compared to labeling the entire dataset, I still don’t want to spend too much [time] on this annotation [work].” Regarding the physical workload, physicians had varying opinions. UD4, who used fewer points for each annotation, felt the physical workload was manageable, while UD5, who meticulously annotated each image, expressed concern: “When there’s a lot of data, this work can get exhausting, both physically and mentally.” engineers also reported reduced time workload (mean = 4.834.834.83) and physical workload (mean = 4.674.674.67). UP1 noted, “Our main role is to guide the training settings and confirm the results, which is a lot easier than having to fix the model whenever something goes wrong, especially when we don’t fully understand the data.
@@ 7. Conclusion
Acknowledgements.
We thank anonymous reviewers for their valuable feedback. This work is supported by grants from the National Natural Science Foundation of China (No. 62372298), the Shanghai Siming Medical Development Foundation (No. SGKY-202405), Shanghai Engineering Research Center of Intelligent Vision and Imaging, Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI), and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (KLIP-HuMaCo).
@@ Appendix B Loading Dataset and Model Supplementary
Atl = Atelectasis, Car = Cardiomegaly, Con = Consolidation, Ede = Edema, Eff = Effusion, Emp = Emphysema, Fib = Fibrosis
Her = Hernia, Inf = Infiltration, Mas = Mass, Nod = Nodule, Ple = Pleural_Thickening, Pne = Pneumonia, Ptx = Pneumothorax
@@ Appendix C Observing and Modifying Attention Supplementary
Figure 10. The comparison of different visualization and interpretability methods is as follows: Refer to caption Original Chest X-ray: Provide the baseline image for analysis. Refer to caption Grad-CAM: Highlights the most influential regions in a coherent, high-level manner. Refer to caption LIME (Top 10 Regions): Delivers fragmented, isolated areas that can be difficult to interpret within a broader diagnostic context. Refer to caption Guided Backpropagation: May introduce noise, potentially obscuring clinically relevant features, particularly in multi-label settings. Refer to caption SHAP: Often results in dispersed or sparse patterns, making interpretation more challenging.
@@ Appendix D Evaluation Metrics Supplementary
[…]
@@ Appendix E User Study Supplementary
OME = Otitis Media with Effusion
EOF = External Auditory Canal Foreign Body
OE = Otitis Externa
EOM = External Auditory Canal Tumor
FOE = Fungal Otitis Externa
CE = Cerumen Impaction
TMP = Tympanic Membrane Perforation